Q1) Generator Produce 25 KW to charge battery and the load is 2.2 KW

1- Calculate the Capacity of the battery ? Please write the Calculation Equation

2- Calculate the specific power of the battery? Please write the Calculation Equation
Answer #1

1)

Generator Voltage : 120 V ( in AC single phase )

Power = 25 kW = V*I

So , I = 25kW / 120 V = 208.33 A

So battery capacity

=Current * time in hours = I * 1 hour = 208.33 Ah

capacity = 208.33 Ah

2)

Specific Power = Energy in kWh / weight

Let weight of battery 4 kg (For lead Acid batteries )

So , Specific Power

= 25kWh/ 4 kg

= 6.25 kWh / kg